# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 4.2 Summary of Field Equations for Each Formulation
$\qquad\hskip 10.0pt+J_{0}g_{mn}\left[-\tfrac{2\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}(K_{bcd}\mathcal{K}^{abcd})-\lambda_{0}^{2}K^{bcd}K_{bcd}+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]$ (4.32)
$\qquad\hskip 10.0pt+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$
$\qquad\hskip 10.0pt+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})~{}~{}~{}.$
In casting the on-shell equations of motion for $g^{mn}$ in this form, starting with either Eq. (2.71) or Eq. (4.19), the following identities prove useful
$Q_{mn}=$${\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+{\bar{\nabla}}_{[m}\text{\textcent}_{n]}-({\rm d}-1){\bar{\nabla}}_{n}\text{\textcent}_{m}+({\rm d}-1)\text{\textcent}_{m}\text{\textcent}_{n}$ (4.33)
${\mathcal{P}}_{mn}=$ ${\bar{{\mathcal{P}}}}_{mn}+{\bar{\nabla}}_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (4.34)
${\mathcal{K}}_{mn}=$ $\hat{R}_{mn}+\tfrac{1}{2}Q_{(mn)}+\tfrac{{\rm d}-1}{2}{\bar{\nabla}}_{(m}\text{\textcent}_{n)}-({\rm d}-1)\text{\textcent}_{m}\text{\textcent}_{n}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}$ (4.35)
$=$ $\hat{R}_{mn}+{\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}~{}~{}~{}.$ (4.36)
@@
We note that in the form above the Einstein Equations for TW gravity are manifestly covariant and projectively invariant. Clearly this is true for the stress energy tensor $\hat{T}_{mn}$, though it may seem at first glance that projective invariance is violated on the left hand side. This is not so, indeed the left-hand side is projectively invariant: the Levi-Civita Ricci tensor $\hat{R}_{mn}$ and Levi-Civita Ricci scalar $\hat{R}$ know not of projective transformations as they depend only on the metric and its derivatives. The projective transformation associated with $R_{ab}$ has been fully hidden in the projectively invariant Palatini field ${\tilde{C}}^{c}{}_{ab}$, leaving $\hat{R}_{mn}$ and $\hat{R}$ projectively invariant. Thus the Einstein equations in the above form are indeed covariant and manifestly projectively invariant.
5 Conclusion
@@
The TW action, built out of the projective Einstein-Hilbert and projective Gauss-Bonnet actions, ties projective geometry to string theory and gravity. The projective connection and the metric are independent in the spirit of the Palatini formalism, giving rise to non-metricity degrees of freedom. The TW field equations then have degrees of freedom from the metric, the diffeomorphism sector, and the Palatini field, which may be encapsulated in different choices of combinations of fields that we have shown are equivalent up to different linear combinations. In this way projective geometry can influence General Relativity by providing sources in the energy-momentum tensor, giving a possible geometric origin to dark energy, dark matter, inflation, and other physical phenomenon.
@@
This paper analyzed pure TW gravity, with no matter or other energy sources: all of the aforementioned applications are of TW gravity itself. Interactions with matter and energy, from standard model sources for instance, are straightforward to include: source terms would appear on the the right hand sides of Eqs. (4.29) rather than being equal to zero. In the form of Einsteins Equations (4.30), interactions with other sources would appear as additions to the stress-energy tensor in the usual way. Indeed, this was already discussed in the previous paper [20]. We look forward to investigating the form of such interactions, in this covariant and projectively invariant formulation, in a future work.
@@
As the equations of motion are different off-shell as derived with respect to different sets of fields, this will have interesting implications at the quantum level in future works. If for instance different gauge fixing procedures produce different results, this could perhaps be related to the trace anomaly in string theory. To iron out such details, the development of a Fadeev-Popov-type of gauge fixing procedure for TW gravity is under consideration. TW gravity itself may be an effective action of a deeper, underlying theory like string theory. Logarithmic running of the coupling constants is already present in the classical theory as demonstrated in this and previous works. Now having a formulation that is both tensorial and projectively invariant will provide a platform for these further investigations.
6 Acknowledgments
@@
The authors would like to thank the Diffeomorphisms and Geometry Research Group and the High Energy Physics Group at the University of Iowa for constructive conversations throughout this work. In particular, V.G.J. Rodgers was the source of much inspiration and elucidation.
@@ Appendix A Conventions and Definitions
Our index convention is as in [20]: Latin indices $a,b,c,\dots$ are over the ${\rm d}$-dimensional manifold $\mathcal{M}$ and Greek indices $\alpha,\beta,\gamma,\dots$ excluding $\lambda$ are over the ${\rm d}+1$-dimensional manifold $\mathcal{N}$, also called the Thomas cone. The connection and curvature tensors on the Thomas cone are succinctly listed in Sections 2.1 and 2.2. The various objects on the Manifold $\mathcal{M}$ are located throughout the paper, so we succinctly list them all here. Parentheses and brackets denote symmetrization and antisymmetrization, respectively. For example, for some arbitrary rank 2 tensor $T$ we define
$T_{(ab)}=T_{ab}+T_{ba}~{}~{}~{},~{}~{}~{}T_{[ab]}=T_{ab}-T_{ba}$ (A.1)
where we do not include a factor of $\frac{1}{2}$.
@@
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}~{}~{}~{}.$ (A.2)
@@
$\alpha_{m}=$ $-\tfrac{1}{d+1}\Gamma^{e}_{\ em}$ (A.3)
$\alpha_{\mu}=$ $(\alpha_{m},\lambda^{-1})$ (A.4)
$g_{m}=$ $-\tfrac{1}{d+1}\partial_{m}\log\sqrt{|g|}=-\tfrac{1}{d+1}\hat{\Gamma}^{e}{}_{me}$ (A.5)
$g_{\mu}=$ $(g_{m},\lambda^{-1})$ (A.6)
¢$\text{\textcent}_{m}=$ $\ g_{m}-\alpha_{m}$ (A.7)
@@
$\bar{\Gamma}^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\tilde{C}^{a}{}_{bc}$ (A.8)
$\Gamma^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+C^{a}{}_{bc}$ (A.9)
$\Pi^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\delta^{a}{}_{(b}g_{c)}+\tilde{C}^{a}{}_{bc}~{}~{}~{}.$ (A.10)
@@
$\hat{\Gamma}^{c}{}_{ab}=$ $\frac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)$ (A.12)
$\tilde{C}^{a}{}_{bc}=$ $C^{a}{}_{bc}-\tfrac{1}{{\rm d}+1}\delta^{a}{}_{(b}C^{m}{}_{c)m}=C^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}~{}~{}~{}.$ (A.13)
@@
$G_{\alpha\beta}$ $=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (A.20)
$G^{\alpha\beta}$ $=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}$ (A.21)
$\Upsilon^{\alpha}=$ $(0,0,\dots,0,\lambda)~{}~{}~{}.$ (A.22)
On the Manifold $\mathcal{M}$, the curvature tensors are
$\hat{R}^{a}{}_{bcd}=$ $\hat{\Gamma}^{a}{}_{b[d,c]}+\hat{\Gamma}^{m}{}_{b[d}\hat{\Gamma}^{a}{}_{c]m}$ (A.23)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{c}C^{a}_{\ bd}-\hat{\nabla}_{d}C^{a}_{\ bc}+C^{a}_{\ ce}C^{e}_{\ bd}-C^{a}_{\ de}C^{e}_{\ bc}$ (A.24)
$\bar{R}^{a}{}_{bcd}=$ $\bar{\Gamma}^{a}{}_{b[d,c]}+\bar{\Gamma}^{m}{}_{b[d}\bar{\Gamma}^{a}{}_{c]m}$ (A.25)
$R^{a}{}_{bcd}=$ $\Gamma^{a}{}_{b[d,c]}+\Gamma^{m}{}_{b[d}\Gamma^{a}{}_{c]m}$ (A.26)
$\qquad$== $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (A.27)
$\mathcal{R}^{a}{}_{bcd}=$ $\Pi^{a}{}_{b[d,c]}+\Pi^{m}{}_{b[d}\Pi^{a}{}_{c]m}$ (A.28)
$\mathcal{K}^{a}{}_{bcd}=$ $R^{a}{}_{bcd}+\delta^{a}_{\ [\,c}{}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (A.29)
$\qquad$== $\mathcal{R}^{a}_{\ bcd}+\delta^{a}_{\ [c}\mathcal{D}_{d]b}$ (A.30)
$=$ $\bar{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}~{}~{}~{}.$ (A.31)
@@
$\mathcal{K}_{bcd}=$$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$
@@
$=$ ¯$\bar{\nabla}_{[c}\bar{\mathcal{P}}_{d]b}~{}~{}~{}.$ (A.32)
Ricci tensors are always constructed by contracting over the first and third indices, and Ricci scalars are the trace of their associated Ricci tensors as seen below
$\mathcal{K}_{bd}=$ $\mathcal{K}^{a}_{\ bad}=\mathcal{R}_{bd}+(d-1)\mathcal{D}_{bd}=R_{bd}+d\mathcal{P}_{db}-\mathcal{P}_{bd}$
$=$ $\bar{R}_{mn}+(d-1)\bar{\mathcal{P}}_{mn}$ (A.33)
$\mathcal{K}=$ $g^{ab}\mathcal{K}_{ab}=\mathcal{R}+(d-1)\mathcal{D}=R+(d-1)\mathcal{P}$
$=$ $\bar{R}+(d-1)\bar{\mathcal{P}}$ (A.34)
$\mathcal{R}_{bd}=$ $\mathcal{R}^{a}_{\ bad}$ (A.35)
$R_{bd}=$ $R^{a}_{\ bad}=\hat{R}^{a}_{\ bad}+Q^{a}_{\ bad}$
$\qquad$== $\hat{R}_{bd}+Q_{bd}$ (A.36)
$\bar{R}_{bd}=$ $\bar{R}^{a}{}_{bad}$
$\qquad$== $\tfrac{1}{2}R_{(bd)}+(d-1)\nabla_{(b}\text{\textcent}_{d)}+(d-1)\text{\textcent}_{b}\text{\textcent}_{d}$ (A.37)
$\mathcal{R}=$ $g^{bd}\mathcal{R}_{bd}$ (A.38)
$R=$ $g^{bd}R_{bd}=g^{bd}\hat{R}_{bd}+g^{bd}Q_{bd}=\hat{R}+Q$ (A.39)
$\bar{R}=$ $g^{bd}\bar{R}_{bd}=R+2(d-1)\nabla^{m}\text{\textcent}_{m}+(d-1)\text{\textcent}_{m}\text{\textcent}^{m}$ (A.40)
$\mathcal{P}=$ $g^{ab}\mathcal{P}_{ab}~{}~{}~{},~{}~{}~{}\mathcal{D}=g^{ab}\mathcal{D}_{ab}~{}~{}~{},~{}~{}~{}Q=g^{ab}Q_{ab}~{}~{}~{}.$ (A.41)
@@
$\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}=\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (A.42)
$\bar{\mathcal{P}}=g^{mn}\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}--g^{mn}\breve{\nabla}_{m}g_{n}-g^{mn}g_{m}g_{n}=\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}~{}~{}.$ (A.43)

### v3

@@ IV.2 Summary of Field Equations for Each Formulation
$\qquad\hskip 6.0pt-g_{mn}\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}(K_{bcd}\mathcal{K}^{abcd})+2\mathbb{L}_{mn}-g_{mn}\mathbb{L}$ (243)
where
$\tilde{Q}_{mn}=$${\bar{\nabla}}_{a}{\tilde{C}}^{a}_{\ mn}+{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{a}_{\ nb}$ (244)
$\tilde{Q}=$${\bar{\nabla}}_{a}{\tilde{C}}^{am}_{\ \ \ m}-{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{am}_{\ \ \ b}~{}.$ (245)
The most straightforward way to arrive at the on-shell form of the stress-energy tensor $\hat{T}_{mn}$ in Eq. (IV.2) is to put $\mathbb{T}_{mn}$ in Eq. (III.3.3) on-shell. That is, set $\mathbb{D}^{ab}$ and $\tilde{\mathbb{E}}_{c}^{\ ab}$ to vanish. Alternatively, to arrive at Eq. (IV.2) from either Eq. (83) or Eq. (228), Eqs. (235), (136), and (138) prove useful.
@@
Recall that in the form above the Einstein equations for TW gravity are covariant and manifestly projective invariant. Clearly this is true for the stress energy tensor $\hat{T}_{mn}$, though it may seem at first glance that projective invariance is violated on the left hand side of the Einstein equations. This is not so, indeed the left-hand side is projectively invariant: the Levi-Civita Ricci tensor $\hat{R}_{mn}$ and Levi-Civita Ricci scalar $\hat{R}$ know not of projective transformations as they depend only on the metric and its derivatives. The projective transformation associated with $R_{ab}$ has been fully hidden in the projectively invariant Palatini field ${\tilde{C}}^{c}{}_{ab}$, leaving $\hat{R}_{mn}$ and $\hat{R}$ projectively invariant. Thus the Einstein equations in the above form are indeed covariant and manifestly projective invariant.
IV.3 Solutions
In this section, we revisit the de Sitter space solution originally presented in [19] in constant volume coordinates $g_{a}=0$. We demonstrate here that the de Sitter space solution is still valid even with non-constant volume coordinates $g_{a}\neq 0$. We start with a brief overview of de Sitter space. A more thorough review can be found in the literature [47, 48, 49, 50] in the east coast metric sign convention (mostly plus) or [51] in the west coast metric sign convention (mostly minus). De Sitter space in ${\rm d}$-dimensions is Minkowski space in ${\rm d}+1$-dimensions constrained on a spacetime surface of constant de Sitter radius $L$
$g_{ab}dx^{a}dx^{b}=$ $\eta_{ab}dx^{a}dx^{b}-dz^{2}$ (246)
$\eta_{ab}x^{a}x^{b}-z^{2}=$ $-L^{2}$ (247)
$\eta_{ab}=$ $\text{diag}(1,-1,-1,\cdots,-1)~{}.$ (248)
Here we are using the west coast metric sign convention as we do throughout the entire paper. Note that the coordinate $z$ is not the projective coordinate $\lambda$. This is simply an extra coordinate typically used to succinctly encode ${\rm d}$-dimensional de Sitter space as above and has nothing to do with the ${\rm d}+1$-dimensional Thomas cone. It is straightforward to remove the $z$-dependence by enforcing the constraint Eq. (247), and then solve for the inverse metric, Christoffel symbols, the Riemann and Ricci curvature tensors, and the Ricci scalar. These results are  [^7] De Sitter space is often described as a space of constant positive spacetime curvature. This physically corresponds to a positive cosmological constant in any convention of which we are aware, though the sign of the Ricci scalar is convention dependent. In our conventions, the Ricci scalar is negative in de Sitter space. The expressions for the Riemann and Ricci tensors each also differ by minus signs across east coast and west coast conventions.
$g_{ab}=$ $\eta_{ab}-\frac{\eta_{ac}\eta_{bd}x^{c}x^{d}}{L^{2}+\eta_{ef}x^{e}x^{f}}~{},~{}g^{ab}=\eta^{ab}+\frac{x^{a}x^{b}}{L^{2}}$ (249)
$\hat{\Gamma}^{c}{}_{ab}=$ $-\frac{x^{c}}{L^{2}}g_{ab}~{},~{}\hat{R}_{abcd}=-\frac{1}{L^{2}}g_{a[c}g_{d]b}$ (250)
$\hat{R}_{ab}=$ $-\frac{{\rm d}-1}{L^{2}}g_{ab}~{},~{}\hat{R}=-\frac{{\rm d}({\rm d}-1)}{L^{2}}~{}.$ (251)
As the Ricci tensor is proportional to the metric, so too is the Einstein tensor
$\hat{G}_{ab}\equiv$ $\hat{R}_{ab}-\frac{1}{2}g_{ab}\hat{R}=\Lambda g_{ab}$ (252)
$\Lambda=$ $\frac{({\rm d}-1)({\rm d}-2)}{2L^{2}}$ (253)
with the proportionality factor $\Lambda$ the cosmological constant. Here we see the relationship between the cosmological constant $\Lambda$ and the de Sitter radius $L$. Note that in ${\rm d}=1,2$ the cosmological constant $\Lambda$ vanishes, and thus so too does the Einstein tensor $\hat{G}_{ab}$.
The solution to the equations of motion of TW-gravity in de Sitter space is
$\tilde{C}^{a}{}_{bc}$ $=0~{}~{}~{},~{}~{}~{}\bar{{\mathcal{P}}}_{ab}=\mu g_{ab}~{}~{}~{},~{}~{}~{}\bar{R}_{abcd}=\hat{R}_{abcd}$ (254)
with $\mu$ an arbitrary constant and $\hat{R}_{abcd}$ as above. The $\Pi^{a}{}_{bc}$ equations of motion are trivially satisfied. For the diffeomorphism field and Einstein equations, we restore the coefficients $J_{0}$ and $\kappa_{0}^{-1}$, respectively, as these were factored out of the variation, Eq. (220)
$0=J_{0}\mathbb{D}^{ab}=({\rm d}-1)\left[({\rm d}-2)({\rm d}-3)J_{0}\left(L^{-2}-\mu\right)-\frac{1}{4\kappa_{0}}\right]g^{ab}$ (255)
$\frac{1}{\kappa_{0}}\hat{G}_{mn}=\left\{\frac{\Lambda_{0}}{\kappa_{0}}+({\rm d}-1)({\rm d}-2)\left[\frac{\mu}{2\kappa_{0}}+({\rm d}-3)({\rm d}-4)J_{0}\left(L^{-2}-\mu\right)^{2}\right]\right\}g_{mn}.$ (256)
For ${\rm d}=1$, the diffeomorphism field equations $\mathbb{D}^{ab}$ are trivially satisfied and the Einstein equations are approximately satisfied in the regime $\Lambda_{0}<<\kappa_{0}$. For ${\rm d}=2,3$, the equations are satisfied only for $\kappa_{0}=\infty$: these models must be purely projective Gauss-Bonnet.
For ${\rm d}\geq 4$, we use Eqs. (252) and (253) to simplify Eqs. (255) and (256) to the forms
$\Lambda=$ $\frac{({\rm d}-1)({\rm d}-2)}{2}\mu+\frac{2}{{\rm d}}({\rm d}-2)\Lambda_{0}$ (257)
$\Lambda_{0}=$ $\frac{{\rm d}({\rm d}-1)}{16({\rm d}-2)({\rm d}-3)J_{0}\kappa_{0}}$ (258)
which for ${\rm d}=4$ become
$\Lambda=3\mu+\Lambda_{0}~{}~{}~{},~{}~{}~{}\Lambda_{0}=\frac{3}{8J_{0}\kappa_{0}}~{}.$ (259)
The solution presented in [19] corresponds to the $\kappa_{0}\to\infty$ limit. As done in that work, a further constraint can be applied to $\Lambda$ upon plugging the solution back into the action, and matching the coefficient of the Einstein-Hilbert term to Newton’s constant. A very recent work [52] investigates perturbative solutions around Minkowski space, $\Lambda_{0}=-3\mu$ in Eq. (259), where perturbations of ${\bar{{\mathcal{P}}}}_{ab}$ act as a form of gravitational radiation.
V Conclusion
In this paper, a covariant and manifestly projective invariant formulation of TW gravity was presented in ${\rm d}$-dimensions. The importance of this formulation is that TW gravity is now written in the way general relativity is typically written, in a covariant form, while being manifestly projective invariant at the same time. A key feature available in this new formulation is the ability to utilize partial integration at the Lagrangian level when deriving field equations, as was demonstrated in this paper. Furthermore, the resulting covariant and manifestly projective invariant field equations are much easier to work with in this form than in their original form.
@@ V Conclusion
The TW action, built out of the projective Einstein-Hilbert and projective Gauss-Bonnet actions, ties projective geometry to string theory and gravity. The projective connection and the metric are independent in the spirit of the Palatini formalism, giving rise to non-metricity degrees of freedom. The TW field equations then have degrees of freedom from the metric, the diffeomorphism sector, and the Palatini field, which may be encapsulated in different combinations of fields that we have shown are equivalent up to linear combinations. In this way projective geometry can influence general relativity by providing sources in the energy-momentum tensor, giving a possible geometric origin to dark energy, dark matter, inflation, and other physical phenomenon.
@@
This paper analyzed pure TW gravity, with no matter or other energy sources: all of the aforementioned applications are of TW gravity itself. Interactions with matter and energy, from standard model sources for instance, are straightforward to include: source terms would appear on the right hand sides of Eqs. (241). In the form of the Einstein equations (242), interactions with other sources would appear as additions to the stress-energy tensor in the usual way. Indeed, this was already discussed in the previous paper [20]. We look forward to investigating the form of such interactions, in this covariant and manifestly projective invariant formulation, in a future work. Indeed, one of the authors (TG) has a work in progress on applications to black hole thermodynamics that considers perturbative solutions around the de Sitter space solution that include traceless parts of ${\bar{{\mathcal{P}}}}_{ab}$.
@@
As the equations of motion are different off-shell as derived with respect to different sets of fields, this will have interesting implications at the quantum level in future works. If for instance different gauge fixing procedures produce different results, this could perhaps be related to the trace anomaly in string theory. To iron out such details, the development of a Fadeev-Popov-type of gauge fixing procedure for TW gravity is under consideration. TW gravity itself may be an effective action of a deeper, underlying theory like string theory. Logarithmic running of the coupling constants is already present in the classical theory as demonstrated in this and previous works. Now having a formulation that is covariant and manifestly projective invariant will provide a platform for these further investigations.
Acknowledgements.
@@
The research of K. S. was supported in part by the endowment of the Ford Foundation Professorship of Physics at Brown University. The authors would like to thank the Diffeomorphisms and Geometry Research Group and the High Energy Physics Group at the University of Iowa for constructive conversations throughout this work. In particular, V.G.J. Rodgers was the source of much inspiration and elucidation.
@@ Appendix A Conventions and Definitions
Through most of the paper, we work in units where $c=1$. Specifically, we do this starting in Sec. II.4. The action we use is that of TW gravity as presented in [20], with addition of cosmological constant as in [24]. Our index convention is as in [20]: Latin indices $a,b,c,\dots$ are over the ${\rm d}$-dimensional manifold $\mathcal{M}$ and Greek indices $\alpha,\beta,\gamma,\dots$ excluding $\lambda$ are over the ${\rm d}+1$-dimensional manifold $\mathcal{N}$, also called the Thomas cone. The connection and curvature tensors on the Thomas cone are succinctly listed in Secs. II.1 and II.2. The various objects on $\mathcal{M}$ are located throughout the paper, so we succinctly list them all in this section. Parentheses and brackets denote symmetrization and antisymmetrization, respectively. For example, for some arbitrary rank 2 tensor $T_{ab}$ we define
$T_{(ab)}=T_{ab}+T_{ba}~{}~{}~{},~{}~{}~{}T_{[ab]}=T_{ab}-T_{ba}$ (260)
where we do not include a factor of $\frac{1}{2}$. We use the west coast metric sign convention, which for diagonal metrics has a positive time component with all spatial components negative. Minkowski space in ${\rm d}=4$ is for instance
$\eta_{mn}=\text{diag}(1,-1,-1,-1)~{}.$ (261)
@@
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}~{}.$ (262)
@@
$\alpha_{m}=$ $-\tfrac{1}{d+1}\Gamma^{e}_{\ em}$ (263)
$\alpha_{\mu}=$ $(\alpha_{m},\lambda^{-1})$ (264)
$g_{m}=$ $-\tfrac{1}{d+1}\partial_{m}\log\sqrt{|g|}=-\tfrac{1}{d+1}\hat{\Gamma}^{e}{}_{me}$ (265)
$g_{\mu}=$ $(g_{m},\lambda^{-1})$ (266)
¢$\text{\textcent}_{m}=$ $\ g_{m}-\alpha_{m}$ (267)
@@
$\bar{\Gamma}^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\tilde{C}^{a}{}_{bc}$ (268)
$\Gamma^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+C^{a}{}_{bc}$ (269)
$\Pi^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\delta^{a}{}_{(b}g_{c)}+\tilde{C}^{a}{}_{bc}~{}.$ (270)
@@
$\hat{\Gamma}^{c}{}_{ab}=$ $\frac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)$ (272)
$\tilde{C}^{a}_{\ bc}=$ $C^{a}_{\ bc}-\tfrac{1}{{\rm d}+1}\delta^{a}_{\ (b}C^{m}_{\ c)m}=C^{a}_{\ bc}-\delta^{a}_{\ (b}\text{\textcent}_{c)}~{}.$ (273)
@@
$G_{\alpha\beta}=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (280)
$G^{\alpha\beta}=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}$ (281)
$\Upsilon^{\alpha}=(0,0,\dots,0,\lambda)~{}.$ (282)
On $\mathcal{M}$, the curvature tensors are
$\hat{R}^{a}{}_{bcd}=$ $\hat{\Gamma}^{a}{}_{b[d,c]}+\hat{\Gamma}^{m}{}_{b[d}\hat{\Gamma}^{a}{}_{c]m}$ (283)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{[c}C^{a}_{\ d]b}-C^{m}_{\ \ b[c}C^{a}_{\ d]m}$ (284)
$\tilde{Q}^{a}_{\ bcd}=$$\hat{\nabla}_{[c}{\tilde{C}}^{a}_{\ d]b}-{\tilde{C}}^{m}_{\ \ b[c}{\tilde{C}}^{a}_{\ d]m}$ (285)
$=$ ¯$\bar{\nabla}_{[c}{\tilde{C}}^{a}_{\ d]b}+{\tilde{C}}^{m}_{\ \ b[c}{\tilde{C}}^{a}_{\ d]m}$ (286)
$\mathcal{R}^{a}{}_{bcd}=$ $\Pi^{a}{}_{b[d,c]}+\Pi^{m}{}_{b[d}\Pi^{a}{}_{c]m}$ (287)
$R^{a}{}_{bcd}=$ $\Gamma^{a}{}_{b[d,c]}+\Gamma^{m}{}_{b[d}\Gamma^{a}{}_{c]m}=\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (288)
$\bar{R}^{a}{}_{bcd}=$ $\bar{\Gamma}^{a}{}_{b[d,c]}+\bar{\Gamma}^{m}{}_{b[d}\bar{\Gamma}^{a}{}_{c]m}=\hat{R}^{a}{}_{bcd}+\tilde{Q}^{a}_{\ bcd}$ (289)
$\mathcal{K}^{a}{}_{bcd}=$ $\mathcal{R}^{a}_{\ bcd}+\delta^{a}_{\ [c}\mathcal{D}_{d]b}$ (290)
$=$ $R^{a}{}_{bcd}+\delta^{a}_{\ [\,c}{}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (291)
$=$ $\bar{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}~{}.$ (292)
@@
$K_{bcd}=$$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$
@@
$=$ ¯$\bar{\nabla}_{[c}\bar{\mathcal{P}}_{d]b}~{}.$ (293)
Ricci tensors are always constructed by contracting over the first and third indices
$\mathcal{R}_{bd}\equiv\mathcal{R}^{a}_{\ bad}$ (294)
$R_{bd}\equiv R^{a}_{\ bad}$ $=\hat{R}_{bd}+Q_{bd}\equiv\hat{R}^{a}_{\ bad}+Q^{a}_{\ bad}$ (295)
$\bar{R}_{bd}\equiv\bar{R}^{a}{}_{bad}$ $=\hat{R}_{bd}+\tilde{Q}_{bd}\equiv\hat{R}^{a}_{\ bad}+\tilde{Q}^{a}_{\ bad}$ (296)
$\mathcal{K}_{bd}\equiv\mathcal{K}^{a}_{\ bad}$ $=\mathcal{R}_{bd}+({\rm d}-1)\mathcal{D}_{bd}$
$=R_{bd}+({\rm d}-1)\mathcal{P}_{bd}-{\rm d}\mathcal{P}_{[bd]}$
$=\bar{R}_{mn}+({\rm d}-1)\bar{\mathcal{P}}_{mn}~{}.$ (297)
Ricci scalars are the trace of their associated Ricci tensors as seen below
$\mathcal{R}\equiv g^{bd}\mathcal{R}_{bd}$ (298)
$R\equiv g^{bd}R_{bd}$ $=\hat{R}+Q\equiv g^{bd}\hat{R}_{bd}+g^{bd}Q_{bd}$ (299)
$\bar{R}\equiv g^{bd}\bar{R}_{bd}$ $=\hat{R}+\tilde{Q}\equiv g^{bd}\hat{R}_{bd}+g^{bd}\tilde{Q}_{bd}$ (300)
$\mathcal{K}\equiv g^{ab}\mathcal{K}_{ab}$ $=\mathcal{R}+({\rm d}-1)\mathcal{D}$ (301)
$=R+({\rm d}-1)\mathcal{P}$
$=\bar{R}+({\rm d}-1)\bar{\mathcal{P}}$ (302)
$\mathcal{P}=g^{ab}\mathcal{P}_{ab}$ $~{},~{}\mathcal{D}=g^{ab}\mathcal{D}_{ab}~{},~{}Q=g^{ab}Q_{ab}~{}.$ (303)
@@
$\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}$ (304)
$=$ $\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (305)
$\bar{\mathcal{P}}=g^{mn}\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}-g^{mn}\breve{\nabla}_{m}g_{n}-g^{mn}g_{m}g_{n}$ (306)
$=$ $\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}.$ (307)